OpenStudy (anonymous):

woudl the slope for 6x-5y=12 be 6/5?

OpenStudy (anonymous):

y = (6/5 ) x - 12/5 yes

OpenStudy (anonymous):

-5y=12-6x y=12/5+6/5x from y=mx+c where m is a gradient,then the gradient of y=12/5+6/5x would be 6/5.

OpenStudy (anonymous):

the point would be (0,-2)